Find the equation of the line passes through the point (3, −2) and is perpendicular to the line 3x − 2y = 4.


An environmental study of a certain community suggests that the average daily level of pollution in the air will be Q(p) = √ 0.6p + 20 units when the population is p thousand. It is estimated that after t years the population will be p(t) = 9 + 0.5t2  thousand.

(a) Express the level of pollution in the air as a function of time.

(b) compute the level of pollution after 5 years from now.

(c) When will the pollution level reach 10 units? 



1
Expert's answer
2021-01-29T12:48:47-0500

Solution 1:

Let the slope of the required line be "m_1"

Given line: "3x - 2y = 4"

"\Rightarrow 3x - 4 = 2y
\\ \Rightarrow y=\dfrac{3x}2-2"

On comparing with "y=mx+c" , we get,

"m=\dfrac32" , i.e. slope of given line.

Since, required line is perpendicular to this given line, product of their slopes is -1.

"m\times m_1=-1
\\ \Rightarrow \dfrac32 \times m_1=-1
\\ \Rightarrow m_1=-1 \times \dfrac23=\dfrac{-2}3"

Also, given point is (3, -2), i.e. "x_1=3,\ y_1=-2"

Now, equation of line with slope and a point is:

"y-y_1=m_1(x-x_1)"

Putting values

"y-(-2)=\dfrac{-2}3(x-3)"

"\Rightarrow y+2=\dfrac{-2x}3+2"

"\Rightarrow y=\dfrac{-2x}3"

"\Rightarrow 2x+3y=0"

Answer: "2x+3y=0" .


Solution 2:

Given, Q(p)"=\sqrt{0.6p+20}"

And p(t)"=9+0.5t^2"

(a): Q(p) is level of pollution and function of p and p(t) is population after t years.

We need to express Q(p) in terms of t, i.e. composition function, "Q \omicron P(t)" .

"Q \omicron P(t)=Q[P(t)]=\sqrt{0.6(9+0.5t^2)+20}"

"=\sqrt{5.4+0.3t^2+20}"

"=\sqrt{25.4+0.3t^2}"

(b): Put t=5

"Q \omicron P(5)=\sqrt{25.4+0.3(5)^2}"

"=\sqrt{25.4+7.5}"

"=\sqrt{32.9}"

"\approx 5.74" units

(c): Now, "Q \omicron P(t)=10" units

"\sqrt{25.4+0.3t^2}=10"

Now solving for t,

"{25.4+0.3t^2}=100" [On squaring both sides]

"\Rightarrow {0.3t^2}=100-25.4=74.6"

"\Rightarrow t^2=\dfrac{74.6}{0.3}\approx248.66"

"\Rightarrow t=\sqrt{248.66}=15.76"

Answer:

(a) "\sqrt{25.4+0.3t^2}"

(b) 5.74 units

(c) 15.76 years.